Use the Binomial Theorem to find the third term in the expression of (n-2p)^6.
Thanks

Answer by venugopalramana(3286)

Use the Binomial Theorem to find the third term in the expression of (n-2p)^6.
Thanks
T(R+1)=(NCR)*(X^R){A^(N-R)}...X=N.....A=-2P.......N=6...R=2(3RD TERM..R+1=3)
T3=(6C2)(N^2)(-2P)^(6-2)=(6*5/2)(N^2(2^4)(P^4)=240*(N^2)(P^4)